The 4-year par value of 1000 bond with mature at 1200 and has annual coupons of 25 earns spot rates s1=3%, s2=2%, s3=1%, and s4=5%. Find the price of the bond and determine its annual yield rate
Answer #1

The Price of Bond (V) = V=\frac{C_i}{(1+s_1)}+\frac{C_2}{(1+s_2)^2}+\frac{C_3}{(1+s_3)^3}+\frac{C_4}{(1+s_4)^4}+\frac{M.V}{(1+s_4)^4}

Year Coupon Spot Rate
1 25 3% 24.27184 25/1.03
2 25 2% 24.02922 25/1.02^2
3 25 1% 24.26475 25/1.01^3
4 25 5% 21 25/1.05^4
4 1200 5% 987 1200/1.05^4
Value of Bond 1080.376 SUM(D3:D7)

Annual Yield of bond investment: required rate of return to make net bond value at zero.

А 1 Year o Cash flow -1080.376 25 25 N w 1225 YTM 4.89% IRR(B2:B6)